Solve using excel

SHOES "R" US bonds are currently selling for $1200. They pay a semi-annual dividend at a stated 7.75% annual coupon rate.

The bond matures in seven and one half years and promises a $1000 par value.
1. What is the semi-annual coupon payment of this bond?
2. What is the yield to maturity for this bond?
Answer #1

1.
Semi annual coupon payment of the bond=7.75%*1000/2=38.75

2.
Yield to maturity=RATE(7.5*2,38.75,-1200,1000)*2=4.570%